ID |
Date |
Author |
Subject |
Text |
 |
239
|
Tue Feb 22 17:32:38 2022 |
David | Current status | Michael is modifying
the settings of the electron cooler -->
changed beam from bunched to coasting |
12x |
238
|
Tue Feb 22 16:09:08 2022 |
David, Yury | Current status | Merger, Tape server (WS3) checks:
DONE, writing ~1 1273kB/s, refreshing
is fine |
12x |
237
|
Tue Feb 22 14:32:04 2022 |
David, Yury | Current status | Merger, Tape server (WS3) checks:
DONE, writing ~1 MB/s, refreshing
is fine |
12x |
236
|
Tue Feb 22 12:23:23 2022 |
CB | Slow comparator thresholds lowered - test run | 13:20 Stopped run.
Slow comp thresh 0x64 -> 0x32
on all FEEs |
|
235
|
Tue Feb 22 12:01:46 2022 |
RSS, OF, ML, CB | Beam OFF, Target ON Run 54 | 01:00
Beam optimisation since 09:25.
Target ON and open to the ring |
9x |
234
|
Tue Feb 22 10:30:00 2022 |
RSS, OF, ML | Beam OFF, Target ON Run 54 | 11:30
Beam optimisation. No beam since
09:25. |
10x |
233
|
Tue Feb 22 10:06:50 2022 |
RSS, OF, ML | PMT channel | YR00DA1ZB is the channel for PMT rate. |
|
232
|
Tue Feb 22 09:10:48 2022 |
OH | Analysis of R54_14 | Results of the fortran analyser programs for
R54_14
|
|
231
|
Tue Feb 22 08:31:25 2022 |
RSS, OF, ML | Beam OFF, Target ON Run 54 | 09:30
Beam optimisation. No beam since
09:25. |
10x |
230
|
Tue Feb 22 07:27:28 2022 |
RSS, OF, ML | Beam ON, Target ON Run 54 | 08:30
Beam ON ~1.0E7 pps
Target ON and open to the ring |
10x |
229
|
Tue Feb 22 06:30:09 2022 |
RSS, OF | Beam ON, Target ON Run 54 | 07:30
Beam ON ~8.0E6 pps
Target ON and open to the ring |
10x |
228
|
Tue Feb 22 05:47:18 2022 |
JM, CB, LV, PM, AK | Beam ON, Target ON Run 54 - 2 events at 10MeV! | 06:47
Beam ON ~2.0E6 pps, still fluctates
between 1e6 and 6e6 (like a step function) |
11x |
227
|
Tue Feb 22 05:12:38 2022 |
JM, CB, LV, PM, AK | Beam ON, Target ON Run 54 | 06:00
Beam ON ~1.5E6 pps, still fluctates
between 1e6 and 6e6 (like a step function) |
11x |
226
|
Tue Feb 22 04:09:28 2022 |
JM, CB, LV, PM, AK | Beam ON, Target ON Run 54 - first event at 10MeV ! | 05:00
Beam ON ~2E6 pps, still fluctates
between 2e6 and 6e6 (like a step function) |
11x |
225
|
Tue Feb 22 03:05:06 2022 |
JM, CB, LV, PM, AK | Beam ON, Target ON Run 54 | 04:00
Beam ON ~5.5E6 pps, but it fluctates
between 2e6 and 6e6 (like a step function) |
9x |
224
|
Tue Feb 22 02:02:11 2022 |
JM, CB, LV, PM, AK | Beam ON, Target ON Run 54 | 03:00
Beam ON ~7E6 pps
Target ON and open to the ring |
10x |
223
|
Tue Feb 22 00:32:35 2022 |
JM, CB, LV, PM, AK | Beam ON, Target ON Run 54 | 01:30
Beam ON ~6E6 pps
Target ON and open to the ring |
 |
222
|
Mon Feb 21 21:04:08 2022 |
SHB, JM, CB, OH (remotely) | Schottky RF view from the beam inside | |
|
221
|
Mon Feb 21 20:56:51 2022 |
SHB, JM, CB, OH (remotely) | DAQ connection for ion current | signals from ICT and CRYRADIO are now connected
to AIDA2 and AIDA5 respectively. ICT is set
to factor 200 and CRYRADIO to 40 for now. |
|
220
|
Mon Feb 21 18:03:58 2022 |
SHB, JM, CB, OH (remotely) | Info: signals on ion intensity for experiment | Currently 4 possibilities to have a look
at the beam, all of them are OK to check
whether there is beam or not, but some of |
|
|